OpenStudy (anonymous):

Could I find the two missing angle measures if I know some of the side lengths of a right triangle?

OpenStudy (anonymous):

Yes you could.

OpenStudy (anonymous):

how?

OpenStudy (campbell_st):

use trigonometry... to find 1 angle... sin, cos or tan. you have 2 angles... find the 3rd, since the triangle is right angled you can use angle sum of a triangle..

OpenStudy (anonymous):

can you show me an example?

OpenStudy (campbell_st):

sure  \[\cos(x) = \frac{12}{13}.... then...... x = \cos^{-1}(\frac{12}{13})\] then you ahve found the 1st angle. the other angle is found by 180 - 90 - x hope this helps

OpenStudy (anonymous):

im still confused :/

OpenStudy (campbell_st):

ok... well do you know about trigonometry..?

OpenStudy (anonymous):

i understand for the most part sin cos tan but im not the best at applying it. I was thinking for this it would be the sec cosec and cotan but idk.....i h8 geometry :/

OpenStudy (campbell_st):

ok... so what is cos equal to in terms of the sides of a triangle... opposite, adjacent and hypotenuse..?

OpenStudy (anonymous):

soh-cah-toa.....cos is adjacent over hypotenuse

OpenStudy (campbell_st):

yep so in the diagram cos(x) = 12/13 how do you find an angle using trig.?

OpenStudy (anonymous):

idk :/

OpenStudy (campbell_st):

ok... heard of arccos..... of cos^-1..?

OpenStudy (anonymous):

you have lost me

OpenStudy (campbell_st):

ok... do you have a calculator handy...?

OpenStudy (anonymous):

yeah but i have to type up an example and explain for my project

OpenStudy (campbell_st):

so here is an online calculator http://web2.0calc.com/ press the buttons 2nd cos then 12/13) and = it will tell you the size of angle x in degrees...

OpenStudy (campbell_st):

well its helping your understanding.

OpenStudy (campbell_st):

what answer do you have for the size on angle x

OpenStudy (anonymous):

.3947

OpenStudy (campbell_st):

nope... try again, 2nd then cos the screen says acos( type in 12/13) then press equals...

OpenStudy (anonymous):

when i do that it is cos-1

OpenStudy (campbell_st):

pressing 2nd cos will display acos.... which is the same as cos^-1 it stands for arccos

OpenStudy (anonymous):

.39479117

OpenStudy (campbell_st):

does you calculator display radians or degrees...?

OpenStudy (anonymous):

idk its the online ti-84+silver edition

OpenStudy (campbell_st):

ok... well I have one of them when I press 2nd cos ( 12/13) I get 22.62 can you press the mode key... and look at the 3rd item in the menu... radian degree make sure degree is highlighted

OpenStudy (campbell_st):

also check the next option is func and not pol

OpenStudy (anonymous):

ok now i got the same answer as you :)

OpenStudy (anonymous):

it was in radian :/

OpenStudy (campbell_st):

ok... so with 2 sides you have found an angle. you know the triangle has a right angle so you just need to find the 3rd angle this can be done by angle sum of a triangle... call the 3rd angle A then 180 - 90 - 22.62 = A so A = 67.38 degrees so really you only need 2 sides in a right triangle to find the 2 angles... hope this makes sense.

OpenStudy (anonymous):

ok THANK YOU SO MUCH :)

OpenStudy (campbell_st):

glad to help.